OpenStudy (anonymous):

v^2=u^2+2as u=6 a=2.55 s=9

OpenStudy (perl):

whats the directions?

OpenStudy (bibby):

solve for u^2+2as and then take the square root of both sides

OpenStudy (anonymous):

36+23.1=59.1

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

36+23.1=59.1 ==> u^2+2as

OpenStudy (bibby):

I got a different number. recalculate 2*a*s

OpenStudy (anonymous):

\(\Huge\color{green}{\text{Okay}}\)

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

\(\Huge\color{red}{\text{By the way thnx for your help}}\)

OpenStudy (bibby):

break it up into parts. a=2.55 s=9 \[\huge 2as = 2*9*2.55\] btw, no problem

OpenStudy (anonymous):

Wait so that is 45.9

OpenStudy (anonymous):

then?

OpenStudy (bibby):

correct

OpenStudy (anonymous):

(◕‿-)

OpenStudy (bibby):

(◑‿~)

OpenStudy (bibby):

sooooo 45.9+36=?

OpenStudy (anonymous):

you want value of v=

OpenStudy (anonymous):

81.9 @bibby @hazem-wardat yeah i need the value of v

OpenStudy (anonymous):

(✿◠‿◠)

OpenStudy (anonymous):

v= 9.049862

OpenStudy (bibby):

Exactly. So we have v^2 = 81.9 To get v alone, you take the square root of both sides

OpenStudy (bibby):

hazem pls (◞≼◉ื≽◟ ;益;◞≼◉ื≽◟)

OpenStudy (anonymous):

(◞≼◉ื≽◟ ;益;◞≼◉ื≽◟)

OpenStudy (anonymous):

should i round to the nearest hundredth

OpenStudy (anonymous):

which is 9.05 or to the nearest tenth will it be 9.0

OpenStudy (anonymous):

v=(( u^2)+2as) ^0.5 v = ( (6^2) + (2*2.55*9))^ 0.5

OpenStudy (anonymous):

?????????????????????

OpenStudy (bibby):

the 0.5 or 1/2 power is a square root

OpenStudy (anonymous):

yes you can say this

OpenStudy (anonymous):

right that the 0.5 or 1/2 power is a square root

OpenStudy (anonymous):

so the answer is 0.5 or 9.05

OpenStudy (bibby):

9.05

OpenStudy (anonymous):

Thnx both of you

OpenStudy (bibby):

np (´◉◞౪◟◉)